# TopoBDA: Towards Bezier Deformable Attention for Road Topology Understanding

### v2

@@ 4.1.2 Metrics
In addition to these metrics, a specialized mAP metric is proposed to evaluate topology reasoning in the graph domain. This metric assesses the connectivity and relationships between centerlines and traffic elements. For an edge (connectivity) to be a true positive, both vertices must be correctly detected according to the Fréchet distance for centerlines and IoU criteria for traffic elements. The topology scores are defined as $\text{TOP}_{ll}$ for centerline topology and $\text{TOP}_{lt}$ for centerline-traffic element topology.
@@
The evaluation framework uses the OpenLane-V2 Score (OLS) as the overall metric, which is calculated as the average of several task-specific metrics: Fréchet-based mAP for centerline prediction ($\text{DET}_{l}$), IoU-based mAP for traffic element prediction ($\text{DET}_{t}$), and the topological relationships metrics between centerlines and traffic elements ($\text{TOP}_{ll}$ and $\text{TOP}_{lt}$). The evaluation metric pipeline has been updated from version 1.0 to 1.1 to address previously identified issues, as outlined in the TopoMLP study [66].
The potential shortcomings of the 1.1 baseline are discussed in the TopoMaskV2 study [21], and a modified baseline is proposed, which is referred to as version 1.1m in this study. While the metric itself remains unchanged, the topology scores are re-mapped such that $P(x)+1\times[P(x)>0.05]$, where $P(x)$ represents the topology score. To ensure fair comparison with existing literature, state-of-the-art comparisons are conducted using V1.1, while all our ablations and hyperparameter evaluations are performed using the V1.1m baseline.
@@ 4.2 Experimental Evaluation
In this section, experimental results are presented. First, instance mask formulation is analysed, followed by a comparison of different attention mechanisms and a conceptual analysis of the auxiliary one-to-many set prediction loss. Next, the performance of sensor fusion and SDMap integration, as well as the efficiency of attention mechanisms in decoder implementations, are evaluated. Finally, the results are compared with state-of-the-art benchmarks on the OpenLane-V2 and OpenLane-V1 datasets.
The supplementary material section provides further insights into our experiments (Section S.6), including a comparative analysis of view transformation methods, a comparison of standard and efficient multi-scale implementations, an analysis of the number of encoder and decoder layers, and an evaluation of the influence of backbone, epochs, and sensor fusion on performance.
@@ 4.2.1 Analysis of Instance Mask Formulation
Table 1 presents the impact of indirect instance mask formulation on TopoBDA with the V1.1m metric baseline. These experiments utilize SwinB as the backbone of the architecture. The results show that incorporating the instance mask auxiliary loss (IMAL) significantly improves performance across various metrics, with a 3.7 points increase in DETl score, and a 3.3 point increase in OLSl (See Section 4.1 for the details of metrics and Eq. (17) for OLSl).
Further enhancement with the mask-L1 mix matcher (ML1M) provides additional gains, with a 3.7-point increase in DETl_ch and a 1.4-point increase in OLSl. The improvements underscore the effectiveness of the instance mask auxiliary loss and the mask-L1 mix matcher and validating their adoption for use in TopoBDA. For detailed information about the implementation of instance mask formulation please refer to Section 3.3.
@@ 4.2.2 Comparison of Different Attention Mechanisms
In Table 2(a), a comparative analysis of various attention mechanisms with the V1.1m metric baseline is presented, utilizing SwinB as the 2D backbone. The results show that deformable attention-based decoders significantly outperform both Standard Attention (SA) and Masked Attention (MA). Specifically, Single-Point Deformable Attention (SPDA) outperforms MA by 1.5 points in OLSl, but, it performs the worst among the deformable attention baselines. MPDA4 outperforms SPDA by 3.3 points in OLSl, highlighting the importance of applying the attention mechanism to different regions around the polyline instead of a single fixed center point. Notably, there is negligible performance difference between 4-point (MPDA4) and 16-point (MPDA16) multi-point deformable attentions with only a 0.1 point improvement in favor of MDPA16 in OLSl. The 4-Point Bezier Deformable Attention (BDA), which employs 4 control points, surpasses both MPDA4 and MPDA16 by 0.7 points in DETl_ch and 0.5 points in OLSl. Despite the limited performance improvement, BDA achieves these results with reduced computational complexity, as it does not require converting Bezier control points to polyline points in each transformer decoder layer using matrix multiplication, unlike MPDA. A comparison of the computational complexity of the attention mechanisms is provided in Section 4.2.5 and further details on SPDA, MPDA, and BDA are provided in Section 3.1.
Table 2: The left table presents a comparison of different attention mechanisms, Standard Attention (SA), Masked Attention (MA), Single Point Deformable Attention (SPDA), 4-point Multi-Point Deformable Attention (MPDA4), 16-point Deformable Attention (MPDA16), and Bezier Deformable Attention (BDA). The right table illustrates the impact of ground truth and query repetition (R) in the auxiliary one-to-many set prediction loss strategy, where R=0 indicates the absence of the auxiliary one-to-many set prediction loss. Both table results are on Subset-A of the Openlane-V2 dataset using the V1.1m Baseline
@@ 4.2.4 Performance Analysis of Sensor Fusion and SDMap Integration
Table 3: Sensor Fusion and SDMap Integration Ablations in Subset-A and Subset-B of OpenLane-V2 with V1.1m Baseline. The configuration with the lidar encoder (SECOND) is marked with a dagger (†) [73].
@@
The camera-only configuration serves as the baseline. In Subset-A, adding lidar significantly improves performance, with OLSl increasing from 44.1 to 52.0. Using the lidar encoder (SECOND) [73] further boosts OLSl to 53.2. Similarly, in Subset-B, integrating lidar with the camera results in substantial performance gains, with OLSl increasing from 49.9 to 61.6 and the use of the lidar encoder further increasing to 62.0.
@@
Figure 7 illustrates the advantages of integrating lidar and SDMap with camera sensors in the BEV domain. The corresponding camera views for these examples are shown in Figure S1 in Section S.7. In Figure 7, inaccurate centerline detections are highlighted with circular regions in the BEV images of specific modalities (camera, lidar, and SDMap) or their combinations. ‘GT’ denotes ground truth, while ‘C’, ‘SD’, and ‘L’ represent the camera, SDMap, and lidar, respectively. In the comparison figures, green centerlines denote the ground truth, while red centerlines represent the predictions, overlaid for easier comparison. Lidar proves advantageous for detecting and localizing long-distance centerlines, while SDMap improves centerline localization and orientation, and is beneficial for occluded and unseen areas. When used together, SDMap and lidar minimizes errors, offering the best best performance.
Overall, these results underscore the significant benefits of multi-modal fusion, particularly the integration of lidar and SDMap, in enhancing road topology understanding.
Figure 7: Visual demonstration in the BEV domain showing the impact of lidar and SDMap additions in Subset-A of the OpenLane-V2 dataset. C, SD, and L represent the camera, SDMap, and lidar, respectively. Green polylines indicate the ground truth and red polylines represent predictions. The circular regions highlight the inaccurate regions compared to other reference BEV images.
@@ 4.2.5 Efficiency Analysis of Attention Mechanisms in Decoder Implementations
Table 4: Decoder runtimes (in ms) for different attention types in Torch and ONNX Runtime implementations.
BDA SPDA MPDA4 MPDA16 SA MA
Torch 17.93 17.93 19.93 22.22 15.54 31.51
ONNX Runtime 9.80 9.75 9.90 10.05 14.05 38.15
The time complexity analysis of the decoder for specific attention types is presented in Table 4. For this analysis, an efficient multi-scale implementation [8], which processes different feature scales successively in each decoder layer in a round-robin fashion, is employed.
In this table, the time complexity of Bezier Deformable Attention (BDA), Single Point Deformable Attention (SPDA), Multi Point Deformable Attention (MPDA), Self Attention (SA), and Masked Attention (MA) are compared in both Torch and ONNX Runtime environments. In Torch, SA has the shortest computation duration, likely due to optimization for dynamic graphs. BDA and SPDA have similar durations, with BDA outperforming MPDA4 and MPDA16 by approximately 2ms and 4.2ms, respectively. In ONNX Runtime, SPDA has the shortest duration, followed closely by BDA with a 0.05ms difference. While the performance gap between BDA and MPDA narrows in ONNX Runtime compared to Torch, BDA consistently surpasses both MPDA4 and MPDA16. MA exhibits the highest computational complexity due to the resize operations of generated masks and foreground checks for proper attention mechanisms. Although BDA does not have the shortest computation duration, it has competitive results and it consistently outperforms other attention mechanisms in terms of main evaluation metrics (Table 2(a)).
@@ 4.2.6 Comparison with the State of the Art Results in OpenLane-V2
A STSU [3] C 12.7 43.0 2.9 19.8 29.3
VectorMapNet [39] C 11.1 41.7 2.7 9.2 24.9
MapTR [34] C 8.3 43.5 2.3 8.3 24.2
TopoNet [27] C 28.6 48.6 10.9 23.9 39.8
TopoMLP [66] C 28.5 49.5 21.7 26.9 44.1
Topo2D [25] C 29.1 50.6 22.3 26.2 44.4
TopoLogic [12] C 29.9 47.2 23.9 25.4 44.1
RoadPainter [48] C 30.7 47.7 22.8 27.2 44.6
TopoFormer [47] C 34.7 48.2 24.1 29.5 46.3
TopoMaskV2 [21] C 34.5 53.8 24.5 35.6 49.4
@@
SMERF [43] C + SD 33.4 48.6 15.4 25.4 42.9
TopoLogic [12] C + SD 34.4 48.3 28.9 28.7 47.5
RoadPainter [48] C + SD 36.9 47.1 29.6 29.5 48.2
@@
B TopoNet [27] C 24.3 55.0 6.7 16.7 36.0
TopoMLP [66] C 25.2 63.1 20.7 20.3 44.7
TopoLogic [12] C 25.9 54.7 21.6 17.9 42.3
TopoFormer [47] C 34.8 58.9 23.2 23.3 47.5
TopoMaskV2 [21] C 41.6 61.1 28.7 26.1 51.8
@@
This section presents a comparative analysis of the TopoBDA architecture against other state-of-the-art methods for road topology understanding. Table 5 illustrates this comparison for Subset-A and Subset-B of the OpenLane-V2 dataset, evaluated using the OpenLane-V2 V1.1 metric baseline.
@@ 4.2.7 Comparison with the State of the Art Results in OpenLane-V1
1.5m PersFormer [5] ResNet-50 52.7 0.307 0.319 0.083 0.117
Anchor3DLane [19] ResNet-50 57.5 0.229 0.243 0.079 0.106
@@
LaneCPP [52] EffNet-B7 60.3 0.264 0.310 0.077 0.117
LATR [46] ResNet-50 61.9 0.219 0.259 0.075 0.104
PVALane [80] ResNet-50 62.7 0.232 0.259 0.092 0.118
@@
0.5m PersFormer [5] ResNet-50 43.2 0.229 0.245 0.078 0.106
DV-3DLane (Camera) [44] ResNet-34 52.9 0.173 0.212 0.069 0.098
LATR [46] ResNet-50 54.0 0.171 0.201 0.072 0.099
@@
The proposed TopoBDA architecture outperforms other methods on the OpenLane-V1 dataset, achieving state-of-the-art results across multiple metrics (Table 6). For this analysis, we utilized the ResNet-50 backbone and processed images at 0.5×0.5$\times$0.5 × of the original resolution.
@@
Although LATR achieves the lowest X-error (near) for the 1.5m distance, the difference is minimal, with TopoBDA trailing by only 5 millimeters. Furthermore, TopoBDA surpasses LATR in the X-error (near) metric for the 0.5m distance, outperforming by 14 millimeters. Notably, despite the hyper-parameters being optimized for centerline detection in the OpenLane-V2 dataset, TopoBDA also achieves the best results for lane divider detection in the OpenLane-V1 dataset. Further optimization of score thresholds and hyper-parameters for this task is expected to improve TopoBDA’s performance on the OpenLane-V1 dataset.

### v3

@@ 4.1.2 Metrics
In addition to these metrics, a specialized mAP metric is proposed to evaluate topology reasoning in the graph domain. This metric assesses the connectivity and relationships between centerlines and traffic elements. For an edge (connectivity) to be a true positive, both vertices must be correctly detected according to the Fréchet distance for centerlines and the IoU criteria for traffic elements. The topology scores are defined as $\text{TOP}_{ll}$ for centerline topology and $\text{TOP}_{lt}$ for centerline-traffic element topology.
@@
The evaluation framework uses the OpenLane-V2 Score (OLS) as the overall metric, which is calculated as the average of several task-specific metrics: Fréchet-based mAP for centerline prediction ($\text{DET}_{l}$), IoU-based mAP for traffic element prediction ($\text{DET}_{t}$), and the topological relationships metrics between centerlines and traffic elements ($\text{TOP}_{ll}$ and $\text{TOP}_{lt}$). The evaluation metric pipeline has been updated from version 1.0 to 1.1 to address previously identified issues, as outlined in the TopoMLP study [wu_topomlp_2024].
The potential shortcomings of the 1.1 baseline are discussed in the TopoMaskV2 study [kalfaoglu_topomaskv2_2024], and a modified baseline is proposed, which is referred to as version 1.1m in this study. While the metric itself remains unchanged, the topology scores are re-mapped such that $P(x)+1\times[P(x)>0.05]$, where $P(x)$ represents the topology score. To ensure fair comparison with existing literature, state-of-the-art comparisons are conducted using V1.1, while all our ablations and hyperparameter evaluations are performed using the V1.1m baseline.
@@
4.2 Implementation Details
For consistency with the TopoMLP study [wu_topomlp_2024], the number of control points is set to four, and the Swin-B and ResNet-50 backbones are employed depending on the experimental setting, also following the TopoMLP design choices. However, as shown in Table S6 in Supplementary Section S.6.5, the best performance is achieved with eight control points, although the performance difference compared to four control points is marginal. Furthermore, various backbone architectures are evaluated within TopoBDA, and the results are reported in Table S7 in Section S.6.6. Among the tested models, ConvNeXt-XXL with CLIP pretraining delivers the highest performance.
Additional implementation details, including optimizer configurations, learning rate schedules, batch size, architectural specifications, and hyperparameter settings, are provided in Section S.5 of the supplementary material. This section also includes backbone configurations, transformer parameters, loss coefficients, and other details necessary for reproducibility.
@@ 4.3 Experimental Evaluation
In this section, the experimental results are presented. The experiments are structured as follows. First, the impact of the instance-mask formulation is analyzed. This is followed by a comparative evaluation of various attention mechanisms and a conceptual analysis of the auxiliary one-to-many set prediction loss. Subsequently, the performance of sensor fusion and SDMap integration, as well as the efficiency of different attention mechanisms within decoder implementations, is assessed. Finally, the results are compared against state-of-the-art benchmarks on the OpenLane-V2 and OpenLane-V1 datasets.
The supplementary material section provides further insights into our experiments (Section S.6), including a comparative analysis of view transformation methods, a comparison of standard and efficient multi-scale implementations, a further efficiency analysis of attention types, the number of encoder and decoder layers, an impact of number of control points, an evaluation of the influence of different backbones, and impact of epochs and multi-modality on performance.
@@ 4.3.1 Analysis of Instance Mask Formulation
Table 1 presents the impact of indirect instance mask formulation on TopoBDA with the V1.1m metric baseline. These experiments utilize SwinB as the backbone of the architecture. The impact of other backbone architectures is shown in Supplementary Table S7 in Section S.6.6. The results show that incorporating the instance mask auxiliary loss (IMAL) significantly improves performance across various metrics, with a 3.7 points increase in DETl score, and a 3.0 point increase in OLSl (See Section 4.1 for the details of metrics and Eq. (17) for OLSl).
Further enhancement with the mask-L1 mix matcher (ML1M) provides additional gains, with a 3.7-point increase in DETl_ch and a 1.4-point increase in OLSl. The improvements underscore the effectiveness of the instance mask auxiliary loss and the mask-L1 mix matcher, and validate their adoption for use in TopoBDA. For detailed information about the implementation of instance mask formulation, please refer to Section 3.3 and Supplementary Section S.7.
@@ 4.3.2 Comparison of Different Attention Mechanisms
In Table 2(a), a comparative analysis of various attention mechanisms with the V1.1m metric baseline is presented, utilizing SwinB as the 2D backbone. The results show that deformable attention-based decoders significantly outperform both Standard Attention (SA) and Masked Attention (MA). The number of control points of all experiments is set to 4 to align with the literature [wu_topomlp_2024] and for its practicality (See Supplementary Table S6 in Section S.6.5).
Specifically, Single-Point Deformable Attention (SPDA) outperforms MA by 1.5 points in OLSl, but it performs the worst among the deformable attention baselines. MPDA4 outperforms SPDA by 3.3 points in OLSl, highlighting the importance of applying the attention mechanism to different regions around the polyline instead of a single fixed center point. Notably, there is a negligible performance difference between 4-point (MPDA4) and 16-point (MPDA16) multi-point deformable attentions, with only a 0.1 point improvement in favor of MPDA16 in OLSl. The 4-Point Bezier Deformable Attention (BDA), which employs 4 control points, surpasses both MPDA4 and MPDA16 by 0.7 points in DETl_ch and 0.5 points in OLSl. Despite the limited performance improvement, BDA achieves these results with reduced computational complexity, as it does not require converting Bezier control points to polyline points in each transformer decoder layer using matrix multiplication, unlike MPDA. A comparison of the runtimes of the attention mechanisms is provided in Table 4 of Section 4.3.5, a more detailed computational complexity analysis is presented in Table S3 of Supplementary Section S.6.3, and theoretical details of SPDA, MPDA, and BDA are explained in Section 3.1.
Table 2: The left table presents a comparison of different attention mechanisms, Standard Attention (SA), Masked Attention (MA), Single Point Deformable Attention (SPDA), 4-point Multi-Point Deformable Attention (MPDA4), 16-point Deformable Attention (MPDA16), and Bezier Deformable Attention (BDA). The right table illustrates the impact of ground truth and query repetition (R) in the auxiliary one-to-many set prediction loss strategy, where R=0 indicates the absence of the auxiliary one-to-many set prediction loss. Both table results are on Subset-A of the Openlane-V2 dataset using the V1.1m Baseline.
@@ 4.3.4 Performance Analysis of Sensor Fusion and SDMap Integration
Table 3: Sensor Fusion and SDMap Integration Ablations in Subset-A and Subset-B of OpenLane-V2 with V1.1m Baseline. The configuration with the lidar encoder (SECOND) is marked with a dagger (†) [yan_second_2018].
@@
The camera-only configuration serves as the baseline. In Subset-A, adding lidar significantly improves performance, with OLSl increasing from 44.1 to 52.0. Using the lidar encoder (SECOND) [yan_second_2018] further boosts OLSl to 53.2. Similarly, in Subset-B, integrating lidar with the camera results in substantial performance gains, with OLSl increasing from 49.9 to 61.6 and the use of the lidar encoder further increasing to 62.0.
@@
Figure 9 illustrates the advantages of integrating lidar and SDMap with camera sensors in the BEV domain. The corresponding camera views for these examples are shown in Supplementary Figure S1 of Section S.9. In Figure 9, inaccurate centerline detections are highlighted with circular regions in the BEV images of specific modalities (camera, lidar, and SDMap) or their combinations. ‘GT’ denotes ground truth, while ‘C’, ‘SD’, and ‘L’ represent the camera, SDMap, and lidar, respectively. In the comparison figures, green centerlines denote the ground truth, while red centerlines represent the predictions, overlaid for easier comparison. Lidar proves advantageous for detecting and localizing long-distance centerlines, while SDMap improves centerline localization and orientation, and is beneficial for occluded and unseen areas. When used together, SDMap and lidar minimize errors, offering the best performance.
Overall, these results underscore the significant benefits of multi-modal fusion, particularly the integration of lidar and SDMap, in enhancing road topology understanding. The benefits of multi-modality in higher epoch regimes are shown in Supplementary Table S8 (Section S.6.7).
Figure 9: Visual demonstration in the BEV domain showing the impact of lidar and SDMap additions in Subset-A of the OpenLane-V2 dataset. C, SD, and L represent the camera, SDMap, and lidar, respectively. Green polylines indicate the ground truth, and red polylines represent predictions. The circular regions highlight the inaccurate regions compared to other reference BEV images.
@@ 4.3.5 Efficiency Analysis of Attention Mechanisms in Decoder Implementations
Table 4: Decoder runtimes (in ms) for different attention types in Torch and ONNX. ONNX* indicates inference without auxiliary mask heads.
Runtime Type BDA SPDA MPDA4 MPDA16 SA MA
Torch 18.47 18.55 20.31 22.73 15.96 24.76
ONNX 8.07 8.19 8.11 8.20 11.03 33.41
ONNX* 4.66 4.70 4.67 4.75 9.50 23.09
The time complexity analysis of the decoder for different attention mechanisms is presented in Table 4. This analysis was conducted using the NVIDIA RTX A6000 GPU within the NVIDIA NGC PyTorch 24.10 container. To ensure a fair comparison, the number of cross attention heads for all attention mechanisms is set to 4—matching the 4 control points of BDA—except for MPDA16, which uses 16 heads. For this analysis, an efficient multi-scale implementation [cheng_masked-attention_2022], which processes different feature scales successively in each decoder layer in a round-robin fashion, is employed. A computational and performance comparison between this efficient strategy and the standard multi-scale implementation is provided in Supplementary Table S2 (Section S.6.2).
In this table, the time complexity of Bezier Deformable Attention (BDA), Single Point Deformable Attention (SPDA), Multi Point Deformable Attention (MPDA), Self Attention (SA), and Masked Attention (MA) is compared in both Torch and ONNX runtimes. ONNX* indicates the ONNX runtime of the model without the auxiliary instance mask head, which will be the ideal case in deployment. Runtimes indicate the duration of the complete 10-layer decoder with the specified attention type.
In the Torch runtime, SA has the shortest computation duration, likely due to optimization for dynamic graphs. BDA and SPDA have similar durations, with BDA outperforming MPDA4 and MPDA16 by approximately 1.8 ms and 4.3ms, respectively. In the ONNX runtime, BDA has the shortest duration. While the runtime gap between BDA and MPDA narrows in the ONNX runtime compared to the Torch runtime, BDA consistently surpasses both MPDA4 and MPDA16. This reduction in runtime is likely attributable to more efficient matrix multiplication operations enabled by the static graph optimization in ONNX. MA exhibits the highest computational complexity due to the resize operations of generated masks and foreground checks for proper attention mechanisms.
The comparable runtime between SPDA and BDA is theoretically expected when the number of heads is equal, as the substitution of reference points across attention heads does not inherently alter the computational complexity of the attention mechanism (See Figure 3). However, the requirement of SPDA for additional reference point prediction makes it slightly more computationally expensive. The benefits of BDA over MPDA arise from its removal of matrix multiplication, which can be especially observed in the Torch runtime (See Figure 4). Although BDA does not have the shortest computation duration in the Torch runtime, it is the best in the ONNX runtime, and it consistently outperforms other attention mechanisms in terms of main evaluation metrics (Table 2(a)).
Supplementary Table S3 in Section S.6.3 further demonstrates the number of flops, memory utilization, and number of parameters of onnx models without auxiliary heads, and it can be seen that the onnx runtimes and number of flops are consistent with each other. This supplementary section also examines the runtimes of varying numbers of encoder and decoder layers within the TopoBDA architecture (Supplementary Table S4 and S5).
@@ 4.3.6 Comparison with the State of the Art Results in OpenLane-V2
A STSU [can_structured_2021] C 12.7 43.0 2.9 19.8 29.3
VectorMapNet [liu_vectormapnet_2023] C 11.1 41.7 2.7 9.2 24.9
MapTR [liao_maptr_2023] C 8.3 43.5 2.3 8.3 24.2
TopoNet [li_graph-based_2023] C 28.6 48.6 10.9 23.9 39.8
TopoMLP [wu_topomlp_2024] C 28.5 49.5 21.7 26.9 44.1
Topo2D [li_enhancing_2024] C 29.1 50.6 22.3 26.2 44.4
TopoLogic [fu_topologic_2024] C 29.9 47.2 23.9 25.4 44.1
RoadPainter [ma_roadpainter_2024] C 30.7 47.7 22.8 27.2 44.6
TopoFormer [lv_t2sg_2024] C 34.7 48.2 24.1 29.5 46.3
TopoMaskV2 [kalfaoglu_topomaskv2_2024] C 34.5 53.8 24.5 35.6 49.4
@@
SMERF [luo_augmenting_2023] C + SD 33.4 48.6 15.4 25.4 42.9
TopoLogic [fu_topologic_2024] C + SD 34.4 48.3 28.9 28.7 47.5
RoadPainter [ma_roadpainter_2024] C + SD 36.9 47.1 29.6 29.5 48.2
@@
B TopoNet [li_graph-based_2023] C 24.3 55.0 6.7 16.7 36.0
TopoMLP [wu_topomlp_2024] C 25.2 63.1 20.7 20.3 44.7
TopoLogic [fu_topologic_2024] C 25.9 54.7 21.6 17.9 42.3
TopoFormer [lv_t2sg_2024] C 34.8 58.9 23.2 23.3 47.5
TopoMaskV2 [kalfaoglu_topomaskv2_2024] C 41.6 61.1 28.7 26.1 51.8
@@
This section presents a comparative analysis of the TopoBDA architecture against other state-of-the-art methods for road topology understanding. Table 5 illustrates this comparison for Subset-A and Subset-B of the OpenLane-V2 dataset, evaluated using the OpenLane-V2 V1.1 metric baseline. To ensure fair comparison with existing literature, input images are downsampled to $0.5\times$ their original resolution prior to processing. The complete details of the image pre-processing pipeline are provided in Supplementary Section S.5.4.
@@ 4.3.7 Comparison with the State of the Art Results in OpenLane-V1
1.5m PersFormer [chen_persformer_2022] ResNet-50 52.7 0.307 0.319 0.083 0.117
Anchor3DLane [huang_anchor3dlane_2023] ResNet-50 57.5 0.229 0.243 0.079 0.106
@@
LaneCPP [pittner_lanecpp_2024] EffNet-B7 60.3 0.264 0.310 0.077 0.117
LATR [luo_latr_2023] ResNet-50 61.9 0.219 0.259 0.075 0.104
PVALane [zheng_pvalane_2024] ResNet-50 62.7 0.232 0.259 0.092 0.118
@@
0.5m PersFormer [chen_persformer_2022] ResNet-50 43.2 0.229 0.245 0.078 0.106
DV-3DLane (Camera) [luo_dv-3dlane_2024] ResNet-34 52.9 0.173 0.212 0.069 0.098
LATR [luo_latr_2023] ResNet-50 54.0 0.171 0.201 0.072 0.099
@@
The proposed TopoBDA architecture outperforms other methods on the OpenLane-V1 dataset, achieving state-of-the-art results across multiple metrics (Table 6). For this analysis, we utilized the ResNet-50 backbone and processed images at $0.5\times$ of the original resolution.
@@
Although LATR achieves the lowest X-error (near) for the 1.5m distance, the difference is minimal, with TopoBDA trailing by only 5 millimeters. Furthermore, TopoBDA surpasses LATR in the X-error (near) metric for the 0.5m distance, outperforming by 14 millimeters. Notably, despite the hyperparameters being optimized for centerline detection in the OpenLane-V2 dataset, TopoBDA also achieves the best results for lane divider detection in the OpenLane-V1 dataset. Further optimization of score thresholds and hyperparameters for this task is expected to improve TopoBDA’s performance on the OpenLane-V1 dataset.